America decides to use income approach to calculate GDP with the following: compensation of employees $4000 proprietors income $1500, corporate profit $1200, rental income $440, net interest $150,and has earnings of foreigners in America $50,earnings of citizens from America in foreign country $50, depreciation $40, subsidies $40 and indirect business taxes $30
Calculate the following
a national income
b GDP at market price
c GNP at factor price
d NDP at market price
1
Expert's answer
2020-03-02T08:21:57-0500

a national income=4000+1500+1200+440+150+50-40=7300

b GDP=7300+40+30-50=7320

c GNP=7320+50-50=7320

d NDP=7320-40-30=7250